Circular No. 5715

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

G272.2–3.2 P. F. Winkler and G. J. Hanson, Middlebury College; and M. M.

Phillips, Cerro Tololo Interamerican Observatory, report: “We have obtained observations which show that the optical nebulosity near the center of G272.2–3.2 (IAUC 5709) has characteristics typical of supernova remnant filaments. Emission-line CCD images from the CTIO 0.6-m Schmidt telescope show a faint filament centered at α = 9^{h}06^{m}40^{s}.4, δ = −52^{o}06′42′′ (equinox J2000.0), ∼ 1′ northwest of the x-ray center. The images suggest that [S II] emission (671.6, 673.1 nm) is > ∼ that in Hα, indicative of shock-heated material. A spectrum of this filament from the CTIO 1.5-m telescope indicates a [S II]/Hα ratio ∼ 1.4, and shows [N II] 658.3-nm and [O II] 732.5-nm emission as well, all typical of SNRs. Other nearby nebulosity includes a small complex of brighter filaments 4′ west (near α = 9^{h}06^{m}12^{s}.0, δ = −52^{o}06′52′′), and a faint filament 7′ north and west (at α = 9^{h}06^{m}22^{s}.0, δ = −52^{o}00′51′′). The image data suggest [SII]/Hα ratios between 0.5 and 1.0 in the latter two filaments, but no spectra have been obtained. No diffuse emission is evident in continuum images of the field.”

PERIODIC COMET SWIFT-TUTTLE (1992t) Extension of the ephemeris on IAUC 5671:

1993 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Mar. 4 22^{h}07.^{m}58 −47^{o}02.^{′}1 2.229 1.634 42.^{o}0 24.^{o}0  9.4
9 22 19.90 −49 01.1 2.223 1.694 46.2 25.0  9.7
14 22 33.09 −51 01.8 2.214 1.754 50.2 25.8  9.9
19 22 47.32 −53 04.3 2.205 1.814 54.3 26.5 10.1
24 23 02.80 −55 08.3 2.195 1.874 58.3 26.9 10.3
29 23 19.75 −57 13.3 2.186 1.935 62.2 27.2 10.5
Apr. 3 23 38.47 −59 18.2 2.179 1.995 66.1 27.3 10.7
8 23 59.30 −61 21.3 2.174 2.055 69.8 27.2 10.9
13 0 22.65 −63 20.6 2.172 2.116 73.4 27.0 11.1
18 0 48.92 −65 13.3 2.174 2.175 76.7 26.7 11.2
23 1 18.48 −66 56.2 2.181 2.235 79.8 26.3 11.4
28 1 51.54 −68 25.2 2.194 2.295 82.7 25.8 11.6
May 3 2 27.95 −69 36.3 2.212 2.354 85.3 25.3 11.8
8 3 07.1 −70 26.0 2.236 2.413 87.5 24.7 12.0
13 3 47.8 −70 51.9 2.267 2.472 89.4 24.1 12.2

1993 February 25 Daniel W. E. Green